# Semantic Edge Computing and Semantic Communications in 6G Networks: A Unifying Survey and Research Challenges

### v2

@@ 4.2.2 Task-Level Coding
Shao et al. [92] utilized the Information Bottleneck (IB) theory to induce sparsity in the channel-encoded semantics. Such sparsity is further used to prune the redundant dimensions of such representation, thus reducing the transmitted data size. However, the transmitted representations in [92] are continuous-valued which require an infinite-resolution constellation design for the adopted modulation scheme [130, 131]. Xie et al. [8] resolved the compatibility issue with digital modulation-based communication systems by proposing a discrete task-oriented JSCC where the encoded representations are discretized using a Gumbel-based vector quantization block.
@@
4.3 Advancement in Test-Time Adaptive Systems
4.3.1 Channel-Adaptive Coding
Image transmission. As the parties in a SemCom system experience a dynamic wireless link, the entire channel model changes over time which requires the encoder and decoder to be updated during test time accordingly. To address this issue, Dörner et al. [137] proposed a two-stage training procedure, where first the entire DNN is trained using a stochastic channel model which resembles the channel behavior as closely as possible. Then after deployment, only the receiver DNN is fine-tuned based on the real channel to speed up the adaptation process. Other work [114, 138, 139] utilized reinforcement learning to adapt the encoder DNN weights to channel variations instead of supervised learning. To this end, Aoudia and Hoydis [139] regard the transmitter as a RL agent interacting with the environment consisting of both the channel and receiver. Therefore, the agent learns how to update the encoder DNN parameters to minimize the loss function provided by the environment. Moreover, techniques based on perturbing the latent [113] are also proposed to obtain an estimate of the instantaneous channel gradients which are then used to backpropagate and update the transmitter DNN gradients. O’Shea et al. [140] proposed to learn a full-scale surrogate DNN to accurately model the distribution representing the channel. As such, GANs are leveraged to make the surrogate DNN mimic the ongoing channel behavior [112]. Similarly, Ye et al. [141] leveraged conditional Generative Adversarial Networks (cGANs) for time-varying channel modeling where the received signal corresponding to the pilot data is used as the conditioning information [111]. Inspired by DNN Test-Time Adaptation (TTA) techniques, Park et al. [115] send multiple pilots over changing channel conditions during the training phase, and utilize the model-agnostic meta-learning (MAML) algorithm [142] to meta-train an effective adaptation rule for the decoder. As such, the decoder DNN can be quickly adapted using the learned rules during test time with minimal modifications.
Retransmission-based techniques are also proposed to improve the reliability of SemCom frameworks which leads to variable-length encoded semantics based on the number of retransmissions. For example, Jiang et al. [145] proposed to combine semantic coding with Reed Solomon channel coding and Automatic Repeat Request (ARQ) which is widely-used in conventional communication systems. More specifically, multiple isomorphic semantic encoders and decoders are used in [145], where after one decoder fails to interpret the correct semantics and send acknowledgment (ACK), the source message is re-encoded via the next encoder. Moreover, the code length at the output of the transformer-based semantic encoder is made adaptive to the input sentence length.
@@ 4.3.1 Channel-Adaptive Coding
[144] JSCC, Transformer, Hybrid ARQ Bits per Word, BLEU Text Transmission Adaptive, SL
@@
[115] Meta-Learning, Decoder TTA BER, BLER Data Transmission Adaptive, Bit-Level
@@ 4.3.2 Importance-Aware Coding
Modern wireless communication systems often leverage Orthogonal Frequency Division Multiplexing (OFDM) and MIMO techniques. In order to incorporate SemCom into advanced wireless technologies, several importance-aware semantic transmission approaches are proposed. Zhou et al. [156] proposed a channel prediction algorithm to estimate the future Channel State Information (CSI) in MIMO systems. Then, the sub-carriers having better predicted channel conditions are allocated to the important semantics. Weng et al. [157] investigated the importance-aware SemCom for both single-user and multi-user MIMO. A neural network-based channel estimation algorithm and a singular value decomposition precoding are used to jointly determine the sub-carrier allocation for separate semantics. Liu et al. [158] studied OFDM systems and leveraged DRL to jointly determine the sub-carrier allocation and adaptive quantization based on the semantic importance.
@@
5 Current Limitation and Future Direction
@@ 5.1 Co-Design of Compression, Partitioning, and Communication
Recent studies have made significant advancements in compression, partitioning, and communication. However, much of the existing research tend to focus on only one of these areas while neglecting the others. We believe that SEC could greatly benefit from a joint optimization approach that integrates compression, partitioning, and communication. For example, Mu et al. [159] proposed a novel optimization framework to jointly compress the semantic and optimize the DNN partitioning, which achieving smaller latency compared to prior studies that solely focuses on partitioning [35] or compression [5]. Also, Abdi et al. [48] explored feature compression and channel-adaptive coding, proposing an adaptive compression method based on pruning that effectively enhances DNN inference latency in dynamic wireless environments. Therefore, a promising research direction in SEC could involve the co-design of compression, partitioning, and communication.
@@ 5.2 Security and Privacy of Semantic Systems
While SEC can achieve real-time DNN inference at the edge, the distributed nature will expose the semantic information to adversaries, raising security concerns in the mobile edge system. For example, an adversary can eavesdrop the semantics and add an adversarial perturbation to compromise the task-oriented performance. Zhang et al. [160] provided a rigorous theoretical analysis for adversarial robustness of latent representations using IB. In addition, a large-scale empirical study is conducted to assess the robustness of multiple semantic compression approaches. However, [160] leveraged existing attack algorithms in conventional adversarial machine learning domain which may not be optimized for attacks on the semantics. In parallel to adversarial robustness, the cooperative manner of DNN inference also opens the door to the privacy adversaries which can infer the original input data from the exposed semantic information [161, 162, 163]. While the existing work only focus on optimizing the computation and communication overhead of SEC, we believe the robust learning and privacy-preserving techniques can be a compelling research direction in collaborative intelligence.
@@ 5.3 Multi-Device Collaborative Inference
Although SEC have been extensively studied, most of the existing literature consider spliting DNNs into two parts executed at the mobile and edge separately. However, we point out that it is worth to extend the distributed inference to a more realistic case where multiple devices with heterogeneous resources aim to perform the same task collaboratively. One of the earliest work [164] in the multi-device collaborative inference proposed to divide DNN into multiple parts dynamically and used a matching algorithm for computation offloading. However, it only demonstrated its effectiveness based on simulation. Disabato et al. [165] deployed the multi-device partitioning on a real IoT system but only for a simple 5-layer CNN architecture. Hu and Li [166] considered to split DNN into multiple parts based on its DAG structure. The optimization problem is formulated as a linear programming problem and an off-the-shelf solver is leveraged to find the solution. Jung and Lee [167] considered the distributed DNN inference as a routing problem and proposed a layered graph model to optimize the data transfer and computing node selection. Li et al. [168] proposed to use DRL to achieve a more fine-grained partitioning for complex DNN architectures where conventional approaches may need an exhaustive search. However, [166, 167, 168] didn’t consider the dynamic networking conditions in the real-world setting.
5.4 Adaptive Semantic Systems
Although there is a large body of research on the development of adaptive SEC and SemCom systems [92, 48, 7], most of the existing work focus on static approaches [127, 129]. These methods suffer from noticeable performance loss when used in realistic settings. More specifically, while adaptation techniques are vastly studied to cope with variations in the input space such as domain shifts [142], limited efforts have been made to address the distribution shifts in the latent space [115]. Such distribution shifts may be caused by intentional and unintentional perturbations in the semantic representations. Intentional perturbations are made by adversaries with the goal of compromising the edge computing systems [160], while unintentional perturbations are mainly due to the dynamic communication environment [143]. In order to tackle the intentional attacks, defense mechanisms such as the detection of Out of Distribution (OOD) latent representations need to be studied. On the other hand, developing adaptation techniques specifically tailored for edge computing systems both with and without the help of sensing feedback is another future research opportunity which can address the unintentional variations. Therefore, we believe dynamic approaches deserve further investigation before edge computing can be employed in real-world applications.
@@ 5.5 Multi-Task Scalability/Multi-Modal Versatility
Existing research in both SEC and SemCom tend to develop semantic systems which can handle only a specific task. For example, SEC systems which are designed to perform image classification use different architectures for their DL-based algorithms than the systems that execute semantic segmentation. However, in real-world applications, the same semantic system should be capable of handling various tasks according to the user’s ongoing needs. On the other hand, semantic systems are often specialized for a particular data modality. As an instance, the majority of SemCom systems developed for image transmission applications are incapable of communicating textual data or suffer from unsatisfactory performance when used for text transmission. Therefore, further studies are required to develop scalable semantic systems which are capable of performing multiple tasks and processing various data modalities with little or no modification.
Recently, transformer models [169, 170] are used both for natural language processing [171, 172, 173, 174] and computer vision tasks [175, 176, 177], and proved to be effective for multi-modal data processing. Although there are some work [24, 178] in SemCom which use BERT [179] and Swin Transformer [177] for text and image transmission, the majority of SemCom systems still focus on CNNs which can handle visual data only [35, 15, 48, 4]. As for multiple tasks, Matsubara et al. [86] investigated hybrid Vision Transformers (ViTs) [180] to perform both image classification and object detection. Thus, there is an increasing interest to investigate multi-task multi-modal SEC and SemCom.
@@ 5.6 Real-World Applications
Current research in SEC and SemCom are limited to computer vision and natural language processing tasks remaining other tasks unexplored. Recent studies applied SEC to other applications which requires a real-time performance and low transmission costs. For example, Bahadori et al. [181] used semantic compression to compress the Wi-Fi beamforming feedback information which achieves up to 80% transmission reduction compared to the conventional matrix decomposition approaches. Liu et al. [182] investigated the model partitioning in vehicular networks to achieve low latency and high reliability in Vehicle-to-Vehicle/Infrastructure communication. Li and Bi [39] optimized the DNN splitting point and resource allocation to minimize the energy consumption of sensors in multi-device CSI sensing. We note that other potential applications, such as healthcare and smart home systems, could benefit from SEC and SemCom, and are worth further exploration.

### v3

@@ 4.2.2 Task-Level Coding
Shao et al. [48] utilized the Information Bottleneck (IB) theory to induce sparsity in the channel-encoded semantics. Such sparsity is further used to prune the redundant dimensions of such representation, thus reducing the transmitted data size. However, the transmitted representations in [48] are continuous-valued which require an infinite-resolution constellation design for the adopted modulation scheme [173, 174]. Xie et al. [8], Abdi et al. [59] resolved the compatibility issue with digital modulation-based communication systems by proposing a discrete task-oriented JSCC where the encoded representations are discretized using a Gumbel-based vector quantization block. As such, their method is the task-oriented equivalent of the semantic-level SemCom system presented in [158].
@@
4.2.3 Research Challenges
Channel Model Inaccuracies. All the existing static SemCom systems use a stochastic model for the channel distortions [38, 39, 141]. However, these models are not accurate enough to capture the real-world effects of a dynamic communication environment. Even though dynamic methods try to address this issue by adapting SemCom systems after deployment [178, 48], they still initialize the semantic encoder and decoder using an analytical channel model used during training. To address this challenge, recent studies propose adopting an additional data-driven model for the channel [146, 147]. However, this approach adds more computational burden to the developed SemCom systems, making them difficult to execute and adapt in real-time, especially in resource-constrained devices.
Task and Modality Dependencies. One of the major challenges regarding the proposed SemCom systems is that they require different DNN models for different tasks and data modalities [54]. Consequently, multiple DNNs need to be stored to process various data modalities, or the DNN has to be updated when the desired task changes. Therefore, a unified SemCom system that supports multi-modal data and multiple tasks needs to be studied. To this date, even though several efforts have been made to develop multi-modal and multi-task SemCom systems [175, 168], the desired set of input and output modalities and tasks cannot be dynamically changed at test time.
Optimization Complexity. Existing studies need to learn SemCom systems in an end-to-end manner by simultaneously sampling over the data and channel distributions. However, the end-to-end design of the SemCom encoder, precoder, postcoder, and decoder that can be adaptive to both the data set and channel realizations results in huge training overhead as the learning and communication objectives may not be in line with each other [179]. To address this, Cai et al. [180] proposed a decoupled design, in which the learning codecs and communication blocks are trained separately. However, their proposed method suffers from a performance drop as they used a surrogate optimization objective instead of the end-to-end learning objective.
@@ 4.2.3 Research Challenges
[167] LLM, Transformer PSNR, SSIM, LPIPS Image Transmission Static, SL
[168] Diffusion, cGAN Accuracy, Compression Multi-Modal Transmission Static, SL
[160, 161] Transformer, Diffusion PSNR, LPIPS, MSE, SSIM Image Transmission Static, SL
@@
[158] Digital Modulation, VAE PSNR, Accuracy Image Transmission, Image Classification Static, SL/TL
[159] JSCC, Digital Modulation PSNR, Accuracy, mAP Image Transmission, Image Classification Static/Adaptive, SL/TL
[162] Contrastive Learning, CNN PSNR, SSIM, Accuracy Image Transmission, Image Classification Static, SL/TL
[181] MIMO, Adaptive Feedback, Knowledge Distillation PSNR, CSI Length Image Transmission Adaptive, SL
@@
[182] JSCC, Transformer, Hybrid Automatic Repeat Request (ARQ) Bits per Word, BLEU Text Transmission Adaptive, SL
@@
[150] Meta-Learning, Decoder Test-Time Adaptation (TTA) BER, BLER Data Transmission Adaptive, Bit-Level
@@
[180, 179] MIMO, Deep Unfolding, Surrogate Design Goal Accuracy, Rate, Latency Multi-View Image Classification Adaptive, TL
4.3 Advancement in Test-Time Adaptive Systems
4.3.1 Channel-Adaptive Coding
Image transmission. As the parties in a SemCom system experience a dynamic wireless link, the entire channel model changes over time which requires the encoder and decoder to be updated during test time accordingly. To address this issue, Dörner et al. [184] proposed a two-stage training procedure, where first the entire DNN is trained using a stochastic channel model which resembles the channel behavior as closely as possible. Then after deployment, only the receiver DNN is fine-tuned based on the real channel to speed up the adaptation process. Other work [149, 185, 186] utilized reinforcement learning to adapt the encoder DNN weights to channel variations instead of supervised learning. To this end, Aoudia and Hoydis [186] regard the transmitter as a RL agent interacting with the environment consisting of both the channel and receiver. Therefore, the agent learns how to update the encoder DNN parameters to minimize the loss function provided by the environment. Moreover, techniques based on perturbing the latent [148] are also proposed to obtain an estimate of the instantaneous channel gradients which are then used to backpropagate and update the transmitter DNN gradients. O’Shea et al. [187] proposed to learn a full-scale surrogate DNN to accurately model the distribution representing the channel. As such, GANs are leveraged to make the surrogate DNN mimic the ongoing channel behavior [147]. Similarly, Ye et al. [188] leveraged cGANs for time-varying channel modeling where the received signal corresponding to the pilot data is used as the conditioning information [146]. Inspired by DNN TTA techniques, Park et al. [150] send multiple pilots over changing channel conditions during the training phase, and utilize the model-agnostic meta-learning (MAML) algorithm [189] to meta-train an effective adaptation rule for the decoder. As such, the decoder DNN can be quickly adapted using the learned rules during test time with minimal modifications. Cai et al. [180] proposed a MIMO SemCom system by adding a MIMO precoder block after the semantic encoder, where the precoder takes the CSI as an input and can be adapted based on the ongoing channel response. As the design objectives of the learning (i.e., semantic encoder and decoder) and communication (i.e., MIMO precoder) modules are different in previous studies, the authors of [180] align such goals by using a unified surrogate measure and separately optimizing these modules. Cai et al. [179] further improved the alignment of these design goals through a second phase of training using the end-to-end learning objective. In addition, an information-theoretic formulation is provided in [179] to back up the proposed training method.
Other approaches focus on making the channel feedback size adaptive with respect to the current channel conditions. For example, the authors in [181] proposed a MIMO SemCom system, called SCAN, with a lightweight performance evaluator module that is integrated into the transmitter and can predict the reconstruction quality of each image. The compression level of the CSI feedback matrix is then adjusted based on the predicted performance.
Retransmission-based techniques are also proposed to improve the reliability of SemCom frameworks which leads to variable-length encoded semantics based on the number of retransmissions. For example, Jiang et al. [183] proposed to combine semantic coding with Reed Solomon channel coding and ARQ which is widely-used in conventional communication systems. More specifically, multiple isomorphic semantic encoders and decoders are used in [183], where after one decoder fails to interpret the correct semantics and send acknowledgment (ACK), the source message is re-encoded via the next encoder. Moreover, the code length at the output of the transformer-based semantic encoder is made adaptive to the input sentence length.
@@ 4.3.2 Importance-Aware Coding
Modern wireless communication systems often leverage Orthogonal Frequency Division Multiplexing (OFDM) and MIMO techniques. In order to incorporate SemCom into advanced wireless technologies, several importance-aware semantic transmission approaches are proposed. Zhou et al. [199] proposed a channel prediction algorithm to estimate the future CSI in MIMO systems. Then, the sub-carriers having better predicted channel conditions are allocated to the important semantics. Weng et al. [200] investigated the importance-aware SemCom for both single-user and multi-user MIMO. A neural network-based channel estimation algorithm and a singular value decomposition precoding are used to jointly determine the sub-carrier allocation for separate semantics. Liu et al. [201] studied OFDM systems and leveraged DRL to jointly determine the sub-carrier allocation and adaptive quantization based on the semantic importance.
@@
4.3.3 Research Challenges
Adaptation Complexity. Even though dynamic SemCom systems are developed to tackle distribution shifts in data and communication environments [179, 150], the proposed adaptation mechanisms are computationally expensive. This leads to an intolerable latency in the proposed SemCom systems. Also, such systems fail to respond quickly after a major test-time change. Recent studies [202] address this challenge by introducing a separate lightweight adaptive module instead of adapting the entire SemCom transmitter and receiver chain.
Changing Objectives. The existing works focus on adapting the SemCom systems only based on the observations from the communication channel and input data Cai et al. [180], O’Shea et al. [147]. However, user requirements, communication constraints, resource availability, and device workload may change during test time. Therefore, SemCom systems should be also adaptable based on high-level commands or waveform objectives.
Security Threats. SemCom systems introduce unique security challenges due to their different design goals compared to traditional communication systems. For example, SemCom systems with adaptive rate control mechanisms are proposed, where the number of transmitted symbols depends on the difficulty of the message Yang and Kim [178], Shao et al. [48]. As such, an adversary can focus on injecting semantic noise at the transmitter to compromise the communication efficiency of SemCom systems, apart from their performance. In addition, the attacker’s objective can be the maximization of the semantic content of the stolen data rather than its quantity. Moreover, stealing the knowledge bases in the SemCom systems that employ them can also be identified as a potential threat. On the other hand, SemCom systems are vulnerable to over-the-air attacks (attacks in the latent space) which damage their performance [203]. Another topic that requires further investigation is rendering adversarial transmitters and receivers unable to communicate with the SemCom parties, where an encryption scheme could be potentially employed to establish a secure SemCom link.
Lack of Higher-Level Networking Functionalities. Studied SemCom systems assume a direct link between the transmitter and receiver Abdi et al. [59]. However, retransmission and routing mechanisms need to be developed so that nodes that are multiple hops apart in a SemCom network can communicate with each other. Moreover, multi-user semantic signal fusion/defusion is another challenging area that should be explored more in depth.
5 Future Research Directions
@@ 5.1 Co-Design of Compression, Partitioning, and Communication
Recent studies have achieved significant advances in compression, partitioning, and communication for distributed DNNs. However, most existing research focuses on optimizing only one of these areas in isolation, neglecting potential synergies between them. This narrow focus is understandable given the inherent complexity of joint optimization, which demands extensive cross-domain expertise spanning both SEC and SemCom, along with comprehensive system design.
As outlined in Section 2.4, SEC and SemCom can substantially benefit from integrated designs that unify optimization approaches across both domains. While limited, some research has successfully demonstrated joint optimization can achieve superior performance compared to single-objective approaches. For instance, Mu et al. [204] proposed a novel optimization framework to jointly compress the semantic and optimize the DNN partitioning, which achieving smaller latency compared to prior studies that solely focuses on partitioning [62] or compression [5]. Also, Abdi et al. [71] explored feature compression and channel-adaptive coding, proposing an adaptive compression method based on pruning that effectively enhances DNN inference latency in dynamic wireless environments. Therefore, a promising research direction could involve the co-design of compression, partitioning, and communication.
@@ 5.2 Security and Privacy of Semantic Systems
While distributing DNNs can achieve real-time performance in resource-constrained environments, the distributed nature also expose the semantic information to adversaries, raising security concerns in these semantic systems. An adversary can eavesdrop the semantics and add an adversarial perturbation to compromise the task-oriented performance. A prior study [203] provided a rigorous theoretical analysis for adversarial robustness of latent representations using IB. In addition, it provided a large-scale empirical study to assess the robustness of multiple semantic compression approaches. However, [203] leveraged existing attack algorithms in conventional adversarial machine learning domain which may not be optimized for attacks on the semantics. The adversarial robustness of SEC and SemCom need a more comprehensive assessment with more advanced attack methods.
In addition, adversarial attacks can compromise not only the end-to-end task performance but also the in-network performance. For example, Zhang et al. [139] revealed that entropy coding methods generating variable-length semantics are vulnerable to both noise and adversarial perturbations in the input space. Specifically, adversaries can degrade network performance by manipulating inputs to increase the entropy of latent representations, thereby exploiting the variable-length nature of the coding scheme.
Moreover, the cooperative manner of DNN inference also creates additional privacy risks. Adversaries can potentially reconstruct original input data from exposed semantic information [205, 206, 207]. This reconstruction capability poses serious privacy concerns, particularly in applications involving sensitive personal data.
Current research in SEC and SemCom has predominantly focused on optimizing computational and communication efficiency while largely overlooking these critical security dimensions. We argue that integrating robust learning techniques and privacy-preserving mechanisms represents a compelling and necessary research direction for advancing secure collaborative intelligence systems.
@@ 5.3 Multi-Device Collaborative Inference
Most of the existing literature in SEC and SemCom only consider the scenario where DNNs are deployed on a single mobile and a single edge device. This assumes the mobile-edge system has enough resource to execute large DNN models. However, this assumption becomes increasingly unrealistic, as individual mobile devices and edge servers often lack the computational capacity required for modern large-scale models, particularly contemporary large language models and vision foundation models that demand extensive computational resources [208, 209]. As such, it is worth to extend the distributed inference to a more realistic case where multiple devices with heterogeneous resources are used to perform the same task collaboratively.
Some literature has investigated multi-device collaborative inference. For instance, Mohammed et al. [210] proposed to divide DNNs into multiple parts and dynamically deploy them using a matching algorithm. Disabato et al. [211] implemented the multi-device partitioning on a real IoT system. Hu and Li [99] considered to split DNN into multiple parts based on its DAG structure. The optimization problem is formulated as a linear programming problem and an off-the-shelf solver is leveraged to find the solution. Jung and Lee [212] considered the distributed DNN inference as a routing problem and proposed a layered graph model to optimize the data transfer and computing node selection. Li et al. [213] proposed to use DRL to achieve a more fine-grained partitioning for complex DNN architectures where conventional approaches may need an exhaustive search.
While these studies demonstrated promising results, they face substantial barriers to real-world deployment. The increasing number of devices and heterogeneous operating conditions make the optimization problem more challenging compared to the two-device partitioning scenario. This becomes more severe when the DNN has a large size. Existing work only demonstrated the effectiveness of their methods through simulation or simple DNN architectures [210, 211]. Moreover, the increased problem complexity requires longer optimization times, making it difficult to quickly adapt to changing wireless environments. [99, 212, 213]. This research gap highlights the necessity for more efficient and practical solutions for multi-device collaborative inference.
@@ 5.4 Multi-Task Scalability/Multi-Modal Versatility
Existing research in both SEC and SemCom tend to develop systems optimized for a single task. For example, Matsubara et al. [6] optimized the DNN architecture and training method xclusively for image classification tasks, while the SemCom framework proposed in [48] employed distinct DNN models for different datasets. These separated systems will creates substantial computational overhead when multiple tasks must be executed simultaneously as they require to run multiple DNNs separately. This presents a critical barrier in deployment on resource-constrained devices. As such, it urges future research to investigate more general frameworks that are optimized for multiple tasks simultaneously and can effectively handle multi-modal data inputs.
While limited, some recent work has successfully demonstrated the feasibility of multi-task and multi-modal learning in SEC and SemCom. For instance, Matsubara et al. [214] demonstrated that the proposed multi-task semantic compression can effectively reduce end-to-end latency by up to 95% and energy consumption by up to 88% in a joint image classification, object detection, and semantic segmentation scenario. Zhang et al. [175] proposed to use a single DNN with dynamic feature selection to handle multi-modal and multi-task applications, achieving comparable performance to task-specific models while using 71% fewer computation. These promising results underscore the potential of unified approaches to address the scalability challenges inherent in current task-specific semantic systems.
@@ 5.5 Real-World Applications
As discussed in Section 2.1, SEC and SemCom play significant roles in achieving edge intelligence in 6G networks. However, current research has predominantly validated their frameworks using computer vision and natural language processing datasets, creating a notable gap between the demonstrated capabilities and the diverse requirements of real-world applications.
Several recent studies have begun exploring more specialized applications. For example, Bahadori et al. [215] used semantic compression to compress the Wi-Fi beamforming feedback information which achieves up to 80% transmission reduction compared to the conventional matrix decomposition approaches. Liu et al. [216] investigated the model partitioning in vehicular networks to achieve low latency and high reliability in Vehicle-to-Vehicle/Infrastructure communication. Li and Bi [52] optimized the DNN splitting point and resource allocation to minimize the energy consumption of sensors in multi-device CSI sensing.
These domain-specific implementations demonstrate the broader applicability of SEC and SemCom beyond traditional computer vision and NLP tasks. Nevertheless, numerous other potential applications remain largely unexplored such as healthcare monitoring, smart home systems, and industrial IoT, all of which could significantly benefit from SEC and SemCom frameworks. Future research should systematically investigate these diverse application domains to fully realize the potential of semantic approaches in 6G networks.